hi. i am very confused with this math.i try to understand, buts it WAY too confusing for me. please help. the three problems are from the book Thinking With Mathematical Models: Linear and inverse Variation.
the three problems are
5x+7=3x+5
7+3x=5x-5
2.5x-8=5x+12
THAnkS soo much! its on page 38 numbers 26-28. and the only thing it says is: Solve each equation. Show the steps in your solutions. I'm really lost!
-please help in idaho

Answer by stanbon(75887)

5x+7=3x+5
Subtract 3x from both sides to get:
2x+7=5
Subtract 7 from both sides to get:
2x=-2
Divide both sides by 2 to get:
x=-1
------------------
7+3x=5x-5
Subtract 3x from both sides to get:
7=2x-5
Add 5 to both sides to get:
2x=12
Divide both sides by 2 to get:
x=6
--------------
2.5x-8=5x+12
Subtract 2.5x from both sides to get:
-8=2.5x+12
Subtract 12 from both sides to get:
-20=2.5x
Divide both sides by 2.5 to get:
x=-8
Cheers,
Stan H.